# Knowledge Graph Reasoning with Self-supervised Reinforcement Learning

### v1

@@ Abstract
Reinforcement learning (RL) is an effective method of finding reasoning pathways in incomplete knowledge graphs (KGs). To overcome the challenges of a large action space, a self-supervised pre-training method is proposed to warm up the policy network before the RL training stage. To alleviate the distributional mismatch issue in general self-supervised RL (SSRL), in our supervised learning (SL) stage, the agent selects actions based on the policy network and learns from generated labels; this self-generation of labels is the intuition behind the name self-supervised. With this training framework, the information density of our SL objective is increased and the agent is prevented from getting stuck with the early rewarded paths. Our self-supervised RL (SSRL) method improves the performance of RL by pairing it with the wide coverage achieved by SL during pretraining, since the breadth of the SL objective makes it infeasible to train an agent with that alone. We show that our SSRL model meets or exceeds current state-of-the-art results on all Hits@k and mean reciprocal rank (MRR) metrics on four large benchmark KG datasets. This SSRL method can be used as a plug-in for any RL architecture for a KGR task. We adopt two RL architectures, i.e., MINERVA and MultiHopKG as our baseline RL models and experimentally show that our SSRL model consistently outperforms both baselines on all of these four KG reasoning tasks. Full code for the paper available at https://github.com/owenonline/Knowledge-Graph-Reasoning-with-Self-supervised-Reinforcement-Learning.
@@
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
Unlike the common applications of RL such as robotics and video games, formulating KGC as an RL problem presents a challenge of large action space. For example, many nodes in the current KG datasets connect to hundreds or thousands of nodes. To improve exploration efficiency, we propose a self-supervised RL approach (SSRL), where the RL network is warmed up in a supervised manner in order to give the agent a lead for queries with large action space where RL alone performs poorly in.
@@ V-D Performance difference on different datasets
Figure 7: Distribution of relations in the training set. The vertical lines correspond to the accuracy of the agent on that relation that increases over the course of training
@@ V-E Comparison with DeepPath
TABLE VI: Comparison of mean MAP and SR (success rate) for DeepPath with its original SL pretraining method and DeepPath with our SL pretraining method for the link prediction.
@@
TABLE VII: Comparison of MAP scores of DeepPath with its original SL pretraining method and DeepPath with our SL FB15K-237 dataset
@@
TABLE VIII: Comparison of MAP scores of DeepPath with its original SL pretraining method and DeepPath with our SL pretraining method for the link prediction task on the NELL-995 dataset

### v2

@@ Abstract
Reinforcement learning (RL) is an effective method of finding reasoning pathways in incomplete knowledge graphs (KGs). To overcome the challenges of a large action space, a self-supervised pre-training method is proposed to warm up the policy network before the RL training stage. To alleviate the distributional mismatch issue in general self-supervised RL (SSRL), in our supervised learning (SL) stage, the agent selects actions based on the policy network and learns from generated labels; this self-generation of labels is the intuition behind the name self-supervised. With this training framework, the information density of our SL objective is increased and the agent is prevented from getting stuck with the early rewarded paths. Our self-supervised RL (SSRL) method improves the performance of RL by pairing it with the wide coverage achieved by SL during pretraining, since the breadth of the SL objective makes it infeasible to train an agent with that alone. We show that our SSRL model meets or exceeds current state-of-the-art results on all Hits@k and mean reciprocal rank (MRR) metrics on four large benchmark KG datasets. This SSRL method can be used as a plug-in for any RL architecture for a KGR task. We adopt two RL architectures, i.e., MINERVA and MultiHopKG as our baseline RL models and experimentally show that our SSRL model consistently outperforms both baselines on all of these four KG reasoning tasks. Code for our method and scripts to run all experiments can be found in our GitHub at https://github.com/owenonline/Knowledge-Graph-Reasoning-with-Self-supervised-Reinforcement-Learning.
@@ I Introduction
Unlike the common applications of RL such as robotics and video games, formulating KGC as an RL problem presents a challenge of large action spaces. For example, many nodes in the current KG datasets connect to hundreds or thousands of nodes. Similar cardinalities often arise in real world scenarios as well, such as constructing a knowledge graph to represent a large medical corpus. In this case, entities could include diseases, symptoms, treatments, and medications, while relations might define interactions such as ’causes’, ’alleviates’, or ’contraindicated with’. Accurately inferring missing links, such as identifying an effective treatment for a rare combination of symptoms, relies heavily on the ability to make unintuitive but correct connections, which heavily relies on efficient exploration of the full breadth of the graph. To improve exploration efficiency, we propose a self-supervised RL approach (SSRL), where the RL network is warmed up in a supervised manner in order to give the agent a lead for queries with large action space where RL alone performs poorly in.
@@ IV-C Training framework
Our framework consists of an SL stage followed by an RL stage, enabling the agent to gain familiarity with the full breadth of the knowledge graph before learning the paths which maximize reward. The agent trains using the SL stage objective for a pre-set number of steps before the model weights are saved and training continues with the RL stage objective for the remaining steps.
@@ V-D Performance difference on different datasets
Figure 7: Distribution of relations in the training set and final correct response percentage. Distribution is plotted on a logarithmic scale. The colors correspond to how much the correct response percentage for the relation increased over the course of training.
@@ V-E Comparison with DeepPath
TABLE VI: Comparison of mean average precision and SR (success rate) for DeepPath with its original SL pretraining method and DeepPath with our SL pretraining method for the link prediction.
@@
TABLE VII: Comparison of mean average precision scores of DeepPath with its original SL pretraining method and DeepPath with our SL FB15K-237 dataset
@@
TABLE VIII: Comparison of mean average precision scores of DeepPath with its original SL pretraining method and DeepPath with our SL pretraining method for the link prediction task on the NELL-995 dataset